OpenStudy (anonymous):

Help with this calculus problem. Find the derivative of y=(tan^-1)(1/2x) thanks!

OpenStudy (anonymous):

is it \[\tan^{-1}(\frac{x}{2})\]

OpenStudy (anonymous):

\[\tan^{-1} (1/2x)\]

OpenStudy (anonymous):

use chain rule. \[\tan^{-1}(\frac{1}{2x})\]?

OpenStudy (anonymous):

\[\frac{d}{dx}\tan^{-1}(x)=\frac{1}{x^2+1}\] and by the chain rule \[\frac{d}{dx}\tan^{-1}(\frac{1}{2x})=\frac{1}{(\frac{1}{2x})^2+1}\times \frac{d}{dx}\frac{1}{2x}\]

OpenStudy (anonymous):

\[\frac{d}{dx}\frac{1}{2x}=-\frac{1}{2x^2}\] so your initial answer is \[\frac{d}{dx}\tan^{-1}(\frac{1}{2x})=\frac{1}{(\frac{1}{2x})^2+1}\times \frac{-1}{2x^2}\] then a ton of algebra to clean this up

OpenStudy (anonymous):

good from here?

OpenStudy (anonymous):

Yeah thanks!